DFB, Inc. expects earnings next year of $5.01 per share, and it plans to pay a $3.42 dividend to shareholders (assume that is one year from now). DFB will retain $1.59 per share of its earnings to reinvest in new projects that have an expected return of 15.5% per year. Suppose DFB will maintain the same dividend payout rate, retention rate, and return on new investments in the future and will not change its number of outstanding shares. Assume next dividend is due in one year a. What growth rate of earnings would you forecast for DFB? b. If DFBs equity cost of capital is 11.1%, what price would you estimate for DFB stock? c. Suppose instead that DFB paid a dividend of $4.42 per share at the end of this year and retained only $O.59 per share in eamings. That is, it chose to pay a higher dividend instead of reinvesting in as many new projects. If DFB maintains this higher payout rate in the future, what stock price would you estimate for the firm now? Should DFB raise i ts dividend? a. What growth rate of earnings would you forecast for DFB? DFBs growth rate of earnings is 1%. (Round to one decimal place.)
Answer #1

Given:

retention ratio, b = $1.59/$5.01 = 0.32

ROE = 15.5%

a) growth rate, g = ROE * b = 0.155 * 0.32 = 0.04919 = 4.96%

b) Price = D1/(r-g) = $3.42/(0.111-0.0496) = $55.70

c) b = $0.59/$5.01 = 0.12; ROE = 0.155; g = 0.155 * 0.12 = 1.86%

Price = $4.42/(0.111-0.0186) = $47.84

Company should not increase the dividend and must reinvest in new positibe NPV projects.